# Inference on many jumps in nonparametric panel regression models

### v2

@@ 2.2 Hypotheses and Estimators
In this section, we introduce the hypotheses and estimators. We would like to conduct some simultaneous inferences on $\gamma_{j}$ for $j\in\left[N\right]$. The importance of studying individual/group-specific threshold effects is discussed in Zimmert and Lechner, (2019).
First, we are interested in testing the existence of the threshold effects. The null and alternative hypotheses are defined as follows:
@@
Additionally, we aim to test the homogeneity of the threshold effects. In this case, the null and alternative hypotheses are:
$H_{0}^{\left(2\right)}:\gamma_{j}=\gamma_{0}\ \text{for some }\gamma_{0}\text{ }\forall\text{ }j\in\left[N\right],\quad$
$\text{and}\quad H_{a}^{\left(2\right)}:\text{there is no constant}\gamma_{0}\text{ such that }\gamma_{j}=\gamma_{0}\text{ }\forall\ j\in\left[N\right].$
Rejecting of $H_{0}^{\left(2\right)}$ suggests the presence of heterogeneous threshold effects, in which case it is useful to extract some meaningful information from the consistent estimates $\hat{\gamma}_{j}$’s of $\gamma_{j}$’s. For example, the quantiles of $\hat{\gamma}_{j}$’s give us information on how the threshold effects spread among individuals. Providing simultaneous confidence intervals for the threshold effects is also tempting. If we attempt to study the asymptotics of $\hat{\gamma}=(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{N})^{\top}$, the standard central limit theorem may fail due to the high dimensionality $N$ when $N\rightarrow\infty$. One of the key theoretical contributions of the paper is to provide a general framework for making uniform inferences on heterogeneous threshold effects.
Let $\mu_{j}(x):=\mathbb{E}(Y_{jt}|X_{jt}=x)=f_{j}(x)+\gamma_{j}\mathbf{1}_{\{x\geq c_{0j}\}}.$ Then the threshold effect for individual/group $j$ (denoted as $\gamma_{j}$) can be identified as the difference between the left and right limits of $\mu_{j}$ at point $c_{0j}:$ $\gamma_{j}=\mu_{j}(c_{0j}+)-\mu_{j}(c_{0j}-),$ where $\mu_{j}(c_{0j}-)=\lim_{x\uparrow c_{0j}}\mu_{j}(x)$ and $\mu_{j}(c_{0j}+)=\lim_{x\downarrow c_{0j}}\mu_{j}(x).$ To estimate $\gamma_{j}$, we need to estimate both $\mu_{j}(c_{0j}+)$ and $\mu_{j}(c_{0j}-).$ To this aim, it is natural to adopt local linear estimators. Let $\partial_{+}\mu_{j}(\cdot)$ (resp. $\partial_{-}\mu_{j}(\cdot)$) be the right (resp. left) derivative of function $\mu_{j}(\cdot)$ and denote the estimators:
@@
$w_{j,b}^{+}(u,v)$ $=\frac{K((v-u)/b_{j})[S_{j2,b}^{+}(u)-(v-u)S_{j1,b}^{+}(u)]}{S_{j2}^{+}(u)S_{j0}^{+}(u)-S_{j1,b}^{+2}(u)}\mathbf{1}_{\{v\geq u\}},$
@@
and $w_{j,b}^{-}(u,v)$ (resp. $S_{jl,b}^{-}(u)$) is the same as $w_{j,b}^{+}(u,v)$ (resp. $S_{jl,b}^{+}(u)$) with $+$ and $\geq$ therein replaced by -- and <,<,< , respectively. We can thus evaluate the threshold effect near the cut-off point and define our local linear estimator of the average threshold effect for individual/group $j$ as
$\hat{\gamma}_{j}=\hat{\mu}_{j}(c_{0j}+)-\hat{\mu}_{j}(c_{0j}-)=\sum_{t=1}^{T}w_{jt,b}^{+}Y_{jt}-\sum_{t=1}^{T}w_{jt,b}^{-}Y_{jt},$ (2.11)
where the weights take the following form $w_{jt,b}^{+}=w_{j,b}^{+}(c_{0j},X_{jt})$ and $w_{jt,b}^{-}=w_{j,b}^{-}(c_{0j},X_{jt})$.
Under $H_{0}^{\left(1\right)}:$ $\gamma_{j}=0$ for all $j\in\left[N\right]$, we are exclusively testing for the overall significance of threshold effects. Since the function $f_{j}(\cdot)$ is continuous, we would expect that
$\sum_{t=1}^{T}w_{jt,b}^{+}f_{j}(X_{jt})\approx f_{j}(c_{0j}+)=f_{j}(c_{0j}-)\approx\sum_{t=1}^{T}w_{jt,b}^{-}f_{j}(X_{jt}).$
Hence by (1.1), the difference between our estimator and the parameter can be approximately expressed in terms of a weighted average of the error sequence,
@@
where recall that $e_{jt}=\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}$ denotes the composite error in (1.1). Under $H_{0}^{\left(1\right)}$, one would expect $\left|\hat{\gamma}_{j}\right|$ to be small for all $j\in\left[N\right]$. This motivates us to consider the test statistic
@@
is the conditional variance of $(Tb_{j})^{1/2}\hat{\gamma}_{j}$ for standardization purposes. For now, we assume $v_{j}^{2}$ is given, in general, it is unknown and has to be replaced by its estimate. We discuss the consistent estimation of $v_{j}^{2}$ in Subsection 2.4. Under the assumptions in Section 3, it can be shown that all asymptotic results based on $v_{j}^{2}$ remain valid when it is replaced by the consistent estimate $\hat{v}_{j}^{2}$.
@@
where $\tilde{v}_{j}^{2}=(1-1/N)^{2}v_{j}^{2}+\sum_{i\neq j}^{N}v_{i}^{2}/N^{2}$. For practical implementation, we introduce a feasible version of $\mathcal{Q}$ in Algorithm 2 in Section C of the supplementary material.
@@ 2.3 The Unknown Threshold Case
We propose to estimate $c_{0j}$ by $\hat{c}_{j}:=\arg\max_{1\leq i\leq K}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/\hat{v}_{j}(c_{i})|$for $j\in\left[N\right],$ where $\hat{v}_{j}(c_{i})$ is a consistent estimator of $v_{j}(c_{i})$ which will be introduced in Algorithm 3 in Section C of the supplementary material.
@@ 2.4 Test Procedure
In this subsection, we present the steps of the test procedures. To proceed, we begin by discussing the estimation of the error variance, $\sigma_{e,j}^{2},$ of $e_{jt}$ when $c_{0j}$’s are known.
@@
    Recall the definition of $\hat{\gamma}_{j}$ as in Equation (2.11). Regress $Y_{jt}-\hat{\gamma}_{j}$ on $X_{jt}$ using the local linear method with kernel $K\left(\cdot\right)$ and bandwidth $b_{j}$.
@@
    $\hat{\sigma}_{e,j}^{2}=\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|\leq b_{j}\}}\hat{e}_{jt}^{2}}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|\leq b_{j}\}}}$ (2.19)
The estimator $\hat{\sigma}_{e,j}^{2}$ is used to form $\hat{v}_{j}^{2}$ in the Algorithms 1 and 2 in the case of known $c_{0j}$, which are applied to conduct simultaneous tests for $H_{0}^{\left(1\right)}$ and $H_{0}^{\left(2\right)}$. These algorithms are supported by the theoretical GA results in Section 3. The following Algorithm 1 is for testing $H_{0}^{(1)}$.
Algorithm 1 Testing for existence of jumps: i.e., $\gamma_{j}=0$ for all $j$.
@@
3:     Estimate the variance $v_{j}^{2}$ of $(Tb_{j})^{1/2}\hat{\gamma}_{j}$ by $\hat{v}_{j}^{2}=Tb_{j}\sum_{t=1}^{T}\left(w_{jt,b}^{+}-w_{jt,b}^{-}\right)^{2}\hat{\sigma}_{e,j}^{2}$, where $\hat{\sigma}_{e,j}^{2}$ is a consistent estimator for the error variance using the observations close to the threshold $c_{0j}$ as in (2.19).
@@
5:Calculate the feasible test statistic ^$\hat{\mathcal{I}}:=\mbox{max}_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}|/\hat{v}_{j}$.
6:Reject the null hypothesis if ^$\hat{\mathcal{I}}>q_{\alpha}$, where $q_{\alpha}$ is the $(1-\alpha)$ quantile of the Gaussian random variable $\max_{1\leq j\leq N}|Z_{j}|$ with $Z_{j}$’s being independent standard normal variables.
Algorithm 2 in Section C in the supplement is used for testing homogeneity $H_{0}^{(2)}$. For the case where the threshold is unknown, the corresponding algorithm is detailed in Algorithm 3 in the same Section.
@@
From (1.1), we see that the reduced-form error $e_{jt}$ inherits the dependency structure in $\epsilon_{jt}$. However, this dependence does not affect the calculation of critical values, even in the case of fairly strong cross-sectional dependence (e.g., a factor structure). This is mainly due to the localized nature of our kernel-type estimators $\hat{\gamma}_{j}$’s in our test statistic. In fact, we can show that under some mild conditions, as long as the joint density of $X_{it}$ and $X_{jt}$ is not degenerate for any $\left(i,j\right)$ pair (see Assumption 3.6 in Section 3), the covariance terms between $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$ will be of lower order, $O(1/T)$, compared to the variance terms which are $O(1/(Tb))$. For a more detailed discussion on the key intuition on the impact of the dependence structure in a simplified setting, we refer to Section F of the Supplement materials. For a discussion on the issue of bandwidth selection, we refer to Remark 7 in the next section.
@@ 3 Main Theorems
In this section, we state the assumptions for the asymptotic analyses and then study the asymptotic properties of the test statistics. Sections 3.2 and 3.3 present theoretical results on the known threshold case, providing theoretical support for the procedures detailed in Algorithms 1 and 2. Section 3.4 addresses the case of unknown thresholds and examines the validity of Algorithm 3.
@@ 3.1 Assumptions
Recall $\mathcal{I}$ in (2.13), we aim to study its asymptotic properties. To do this, we require a time series version of a non-centered high-dimensional Gaussian approximation. To account for dependence along both the $j$ and $t$ dimensions, we need to impose certain assumptions.
As mentioned above, we will postulate a general spatial and temporal dependence assumption. As for the covariate variable $X_{jt}$ and the confounding variable $U_{jt}$, we assume that they take the form
@@
where $\left\{e_{t}\right\}_{t\in\mathbb{Z}}$ is a sequence of i.i.d. random innovations, $\left\{H_{j}\right\}_{j\in[N]}$ are functions so that $X_{jt}$ are well defined, the innovations $\left\{\tilde{e}_{t}\right\}_{t\in\mathbb{Z}}$ are i.i.d. and independent of $\left\{e_{t}\right\}_{t\in\mathbb{Z}}$, and $\{\tilde{H}_{j}\}_{j\in[N]}$ are measurable functions. Denote $\mathcal{F}_{t}=(e_{t},e_{t-1},\ldots)$ and ~$\tilde{\mathcal{F}}_{t}=(\tilde{e}_{t},\tilde{e}_{t-1},\ldots)$. For the innovation part, the dependence among $\epsilon_{t}:=(\epsilon_{1t},\ldots,\epsilon_{Nt})^{\top}$is allowed to be strong along the cross-sectional dimension (e.g., with a factor structure) and it has to be weak along the time dimension, as we only exploit the sample size in the time series dimension for the estimation of heterogeneous threshold effects. Specifically, we assume that $\epsilon_{t}$ follows an MA($\infty$) process as follows:
@@
Note that we allow for the dependence between $X_{jt}$ and $U_{jt}$. Equation (3.1) is a fairly general setup, where almost all kinds of dependence structures between $X_{jt}$ and $U_{jt}$ can be fulfilled for the following simple reasoning. For any two continuous random variables $X$ and $U$ with the conditional cumulative distribution function of $U$ given $X$ as $F_{U|X}(\cdot|X)$. Let $\tilde{e}=F_{U|X}(U|X)$. Then by Lemma F.1 in Appendix F in Rio, (2017), $\tilde{e}$ is a uniformly distributed random variable on $[0,1]$ and is independent of $X$. Moreover, $U=F_{U|X}^{-1}(\tilde{e}|X).$ It is worth noting that our model and assumption can be extended to multivariate latent variable $U_{jt}$ without introducing additional theoretical insights.
Let $g_{j}(\cdot)$ denote the density function of $X_{jt}.$ For $t\geq s+1,$ let $g_{j,t}(x|\mathcal{F}_{s}):=d\mathbb{P}(X_{jt}\leq x|\mathcal{F}_{s})/dx.$ For any constant $r>0$, we denote
$v_{j,r}(x)=\mathbb{E}(|\varepsilon_{j}(X_{jt},U_{jt})|^{r}|X_{jt}=x).$ (3.3)
We first impose some conditions on the elements of $\eta_{t}.$
@@
Remark 3.
Assumption 3.1 imposes moment conditions on the underlying noise terms. We allow for two possibilities, $(i)$ only some finite order of moments exist for $\eta_{11}$, and $(ii)$ $\eta_{1}$ is sub-exponential. We shall provide conclusions for different sets of moment assumptions.            $\Box$
The following assumption imposes some conditions on the temporal dependence for the processes $(X_{jt})_{t\in\mathbb{Z}}$ and $(\epsilon_{jt})_{t\in\mathbb{Z}}.$ Denote $\mathcal{C}_{j}$ as a set of search values for the time series $j$ containing at least the threshold values $c_{0j}$. If the true threshold is given, then $\mathcal{C}_{j}=\{c_{0j}\}$ is a single point in subsections 3.2 and 3.3. We will specify $\mathcal{C}_{j}$ later in the theorems for the unknown case.
@@
Remark 4.
Assumption 3.2 $(i)$-$(ii)$ essentially assumes weakly temporal dependence for the processes $(X_{jt})_{t\in\mathbb{Z}}$ and $(\epsilon_{jt})_{t\in\mathbb{Z}}$ for any $j\in\left[N\right],$ respectively. No weak dependence conditions are imposed on the cross-sectional dimension. Therefore we can allow for strong cross-sectional dependence over the $j$ dimension. We shall also note that $\beta$ indicates the strength of the dependency structure of the error processes. In the case of $\beta\rightarrow\infty$, the error process becomes temporally independent.            $\Box$

### v3

@@ 2.2 Hypotheses and Estimators
In this section, we introduce the hypotheses and estimators. We would like to conduct simultaneous inferences on $\gamma_{j}$ and $\gamma_{j}^{[1]}$ for $j\in\left[N\right]$. The importance of studying individual/group-specific threshold effects is discussed in Zimmert and Lechner, (2019).
First, we are interested in testing the existence of the threshold effects ($\gamma_{j}$ or $\gamma_{j}^{[1]}$). The null and alternative hypotheses are defined as follows:
@@
The motivation for considering such a test is crucial. In Figure 1, we present a simple comparison with the pooled threshold test in the literature, for example in Fong et al., (2017). It is clear that the conventional pooled tests exhibit very low power due to signal cancellation, with performance only comparable in cases of strictly positive signals.
Figure 1: Empirical power comparison between our uniform testing procedure (black) for the existence of threshold effects and the test based on the pooled linear threshold model (red) at $5\%$ significance level for different proportions of non-zero coefficients ($\gamma_{j}$) on the x-axis. The non-zero $\gamma_{j}$s are drawn from a standard normal distribution (left panel) and an exponential distribution (right panel). The true process is linear, $f_{j}(x)=0.2x$ (for details see DGP7 in Section E of the supplement.)
When $N$ is large, assuming homogeneous threshold effects $\gamma_{j}$ across all $j$ is restrictive, and inferences based on this assumption can be misleading if the threshold effects are heterogeneous. Therefore, it is important to test for heterogeneous threshold effects. In this case, the null and alternative hypotheses are:
$H_{0}^{\left(2\right)}:\gamma_{j}=\gamma_{0}\ \text{for some }\gamma_{0},\text{ }\forall\text{ }j\in\left[N\right],\quad$
and $H_{a}^{\left(2\right)}:\text{there is no constant }\gamma_{0}\text{ such that }\gamma_{j}=\gamma_{0},\text{ }\forall\ j\in\left[N\right].$
We can similarly define $H_{0}^{(1)[1]}$ and $H_{0}^{(2)[1]}$ for the existence and homogeneity tests of the derivative, respectively. Rejecting of $H_{0}^{\left(2\right)}$ suggests the presence of heterogeneous threshold effects. Moreover, providing simultaneous confidence intervals for the threshold effects is also tempting. If we would like to study the asymptotics of $\hat{\gamma}=(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{N})^{\top}$, the standard central limit theorem may fail due to the high dimensionality $N$. One of the key theoretical contributions is to provide a general framework for making uniform inferences on heterogeneous threshold effects.
To estimate $\gamma_{j}$, we need to estimate both $\mu_{j}(c_{0j}+)$ and $\mu_{j}(c_{0j}-).$ To this aim, it is natural to adopt local linear estimators, see for example Fan, (2018). Denote the estimators:
@@
$w_{j,b}^{+}(u,v)$ $=\frac{K((v-u)/b_{j})[S_{j2,b}^{+}(u)-(v-u)S_{j1,b}^{+}(u)]}{S_{j2}^{+}(u)S_{j0}^{+}(u)-S_{j1,b}^{+2}(u)}\mathbf{1}_{\{v\geq u\}},$ (2.8)
$w_{j,b}^{+[1]}(u,v)$ $=\frac{K((v-u)/b_{j})[S_{j1,b}^{+}(u)-(v-u)S_{j0,b}^{+}(u)]}{-S_{j2}^{+}(u)S_{j0}^{+}(u)+S_{j1,b}^{+2}(u)}\mathbf{1}_{\{v\geq u\}}.$ (2.9)
where
@@
and $w_{j,b}^{-}(u,v)$ (resp. $S_{jl,b}^{-}(u)$, $w_{j,b}^{-[1]}(u,v)$) is the same as $w_{j,b}^{+}(u,v)$ (resp. $S_{jl,b}^{+}(u)$, $w_{j,b}^{+[1]}(u,v)$) with $+$ and $\geq$ therein replaced by -- and <,<,< , respectively.
We can thus evaluate the threshold effect near the cut-off point and define our local linear estimator of the average threshold effect for individual/group $j$ as
$\hat{\mu}_{j}(c_{0j}+)=\sum_{t=1}^{T}w_{jt,b}^{+}Y_{jt}\quad\textrm{and}\quad{\partial_{+}\hat{\mu}_{j}}(c_{0j})=\sum_{t=1}^{T}w_{jt,b}^{+[1]}Y_{jt},$ (2.11)
where the weights take the following form $w_{jt,b}^{+}=w_{j,b}^{+}(c_{0j},X_{jt})$ and $w_{jt,b}^{-}=w_{j,b}^{-}(c_{0j},X_{jt})$. Thus we have,
$\hat{\gamma}_{j}=\sum_{t=1}^{T}w_{jt,b}^{+}Y_{jt}-\sum_{t=1}^{T}w_{jt,b}^{-}Y_{jt}\quad\mbox{and}\quad\hat{\gamma}_{j}^{[1]}=\sum_{t=1}^{T}w_{jt,b}^{+[1]}Y_{jt}-\sum_{t=1}^{T}w_{jt,b}^{-[1]}Y_{jt}.$
We shall focus on $\hat{\gamma}_{j}$ from now on, and the algorithms and theorems concerning the statistical properties of $\hat{\gamma}_{j}^{[1]}$ are developed in Section A.5 in the Appendix.
Under $H_{0}^{\left(1\right)}:$ $\gamma_{j}=0$ for all $j\in\left[N\right]$, we are exclusively testing for the overall significance of threshold effects. Since the functions $\tilde{h}_{j}(\cdot)$ and $\tilde{\tau}_{j}(\cdot)$ are continuous, by (2.1), the difference between our estimator and the parameter can be approximately expressed in terms of a weighted average of the error sequence,
@@
where $e_{jt}$ is the composite error in (2.1). Under $H_{0}^{\left(1\right)}$, one would expect $\left|\hat{\gamma}_{j}\right|$ to be small for all $j\in\left[N\right]$. This motivates us to consider the test statistic
@@
is the conditional variance of $(Tb_{j})^{1/2}\hat{\gamma}_{j}$ for standardization purposes. For now, we assume $v_{j}^{2}$ is given, in general, it is unknown and has to be replaced by its estimate. We discuss a consistent estimator of $v_{j}^{2}$ in Subsection 2.4. Under the assumptions in Section 3, it can be shown that all asymptotic results based on $v_{j}^{2}$ remain valid when it is replaced by the consistent estimate $\hat{v}_{j}^{2}$.
@@
where $\tilde{v}_{j}^{2}=(1-1/N)^{2}v_{j}^{2}+\sum_{i\neq j}^{N}v_{i}^{2}/N^{2}$. For practical implementation, we introduce a feasible version of $\mathcal{Q}$ in Algorithm SA-2 in Section D of the Supplementary Material.
@@ 2.3 The Unknown Threshold Case
We propose to estimate $c_{0j}$ by $\hat{c}_{j}:=\arg\max_{1\leq i\leq K}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/\hat{v}_{j}(c_{i})|$for $j\in\left[N\right],$ where $\hat{v}_{j}(c_{i})$ is a consistent estimator of $v_{j}(c_{i})$ which will be introduced in Algorithm SA-3 in Section D of the Supplementary Material.
@@ 2.4 Test Procedure
In this subsection, we present the steps of the test procedures. To proceed, we begin by discussing the estimation of the error variance $\sigma_{e,j}^{2}$ of $e_{jt}$ when $c_{0j}$’s are known.
@@
    Recall the definition of $\hat{\gamma}_{j}$ as in Equation (2.2). Regress $Y_{jt}-\hat{\gamma}_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$ on $X_{jt}$ using the local linear method with kernel $K\left(\cdot\right)$ and bandwidth $b_{j}$.
@@
    $\hat{\sigma}_{e,j}^{2}=\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|\leq b_{j}\}}\hat{e}_{jt}^{2}}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|\leq b_{j}\}}}.$ (2.18)
The estimator $\hat{\sigma}_{e,j}^{2}$ is used to form $\hat{v}_{j}^{2}$ in the Algorithms 1 and SA-2 in the case of known $c_{0j}$, which are applied to conduct simultaneous tests for $H_{0}^{\left(1\right)}$ and $H_{0}^{\left(2\right)}$. These algorithms are supported by the theoretical GA results in Section 3 and Section A in Supplementary Material. The following Algorithm 1 is for testing $H_{0}^{(1)}$.
Algorithm 1 Testing for existence of jumps in conditional means (in slope): i.e., $\gamma_{j}=0$ ( $\gamma_{j}^{[1]}=0$) for all $j$.
@@
3:     Estimate the variance $v_{j}^{2}$ of $(Tb_{j}^{3})^{1/2}\hat{\gamma}_{j}$ by $\hat{v}_{j}^{2}=Tb_{j}\sum_{t=1}^{T}\left(w_{jt,b}^{+}-w_{jt,b}^{-}\right)^{2}\hat{\sigma}_{e,j}^{2}$ ( $\hat{v}_{j}^{2[1]}=Tb_{j}^{3}\sum_{t=1}^{T}\left(w_{jt,b}^{+[1]}-w_{jt,b}^{-[1]}\right)^{2}\hat{\sigma}_{e,j}^{2}$) , where $\hat{\sigma}_{e,j}^{2}$ is a consistent estimator for the error variance using the observations close to the threshold $c_{0j}$ as in (2.18).
@@
5:Calculate the feasible test statistic ^$\hat{\mathcal{I}}:=\mbox{max}_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}|/\hat{v}_{j}$ ( ^$\hat{\mathcal{I}}^{[1]}:=\mbox{max}_{1\leq j\leq N}(Tb_{j}^{3})^{1/2}|\hat{\gamma}_{j}^{[1]}|/\hat{v}_{j}^{[1]}$) .
6:Reject the null hypothesis if ^$\hat{\mathcal{I}}>q_{\alpha}$ ( ^$\hat{\mathcal{I}}^{[1]}>q_{\alpha}$), where $q_{\alpha}$ is the $(1-\alpha)$ quantile of the Gaussian random variable $\max_{1\leq j\leq N}|Z_{j}|$ with $Z_{j}$’s being independent standard normal variables.
@@
Given the well-documented cross-sectional and serial dependence in high dimensional data, it is crucial to develop tests that account for both types of dependence. It is surprising that, according to the above algorithm, $q_{\alpha}$ can be obtained by simulating only i.i.d. Gaussian random variables, without considering the complicated dependency structure of $e_{jt}$. Recall from (2.1), we see that the reduced-form error $e_{jt}$ inherits nontrivial dependency structure in $\epsilon_{jt}$. However, this dependence does not affect the calculation of critical values, even in the case of fairly strong cross-sectional dependence (e.g., a factor structure). This is mainly due to the localized nature of our kernel-type estimators $\hat{\gamma}_{j}$’s in our test statistic. In fact, we can show that under some mild conditions, as long as the joint density of $X_{it}$ and $X_{jt}$ is not degenerate for any $\left(i,j\right)$ pair (see Assumption 3.3 in Section 3), the covariance terms between $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$ will be of lower order compared to the variance terms, $O(1/T)$. For a more detailed discussion on the key intuition on the impact of the dependence structure in a simplified setting, we refer to Section G.1 of the Supplementary Material. For a discussion on the issue of bandwidth selection, we refer to Remark 4 in Section A.2 in the Supplementary Material.
@@ 3 Main Theorems
In this section, we present our main results, with a particular focus on Theorem 3.1, which provide Gaussian approximation results that determine the critical values for testing the existence of threshold effects. Theorem 3.2 extends the result to unknown threshold case and Theorem 3.3 shows the consistency of the estimated threshold locations. Theorem A.1 shows GA for testing homogeneity. Theorems A.2 and A.4 derive the consistency of variance to ensure the validity of our test statistics. Theorem A.3 analyzes the power of the tests for both the threshold effects and homogeneity. Results regarding the derivative cases are presented in Subsection A.5.
@@ 3.1 Assumptions
In this subsection, we state the assumptions for the asymptotic analyses in Section 3.1 and Section A.1 in the Supplementary Material. First of all, we will pose a general spatial and temporal dependence assumption on the data generating processes. As for the covariate variable $X_{jt}$ and the confounding variable $U_{jt}\in\mathbb{R}^{d}$, we assume that they take the form
@@
where $\tilde{H}_{j}=(\tilde{H}_{j1},\ldots,\tilde{H}_{jd})^{\top}$, $\{\tilde{H}_{ji}\}_{j\in[N],i\in[d]}$ and $\left\{H_{j}\right\}_{j\in[N]}$ are functions so that $X_{jt}$ and $U_{jt}$ are well defined. For $t\in\mathbb{Z}$, $e_{t}\in\mathbb{R}$ denote i.i.d. innovations, and $\tilde{e}_{t}\in\mathbb{R}^{\tilde{d}}$ are i.i.d. random vectors with some constant $\tilde{d}>0$, independent of $\left\{e_{t}\right\}_{t\in\mathbb{Z}}$. Denote
$\mathcal{F}_{t}=(e_{t},e_{t-1},\ldots)\quad\textrm{and}\quad\tilde{\mathcal{F}}_{t}=(\tilde{e}_{t},\tilde{e}_{t-1},\ldots).$ (3.2)
For the innovation part $\epsilon_{t}=(\epsilon_{t1},\ldots,\epsilon_{tN})^{\top}$ as in (1.1), the dependence among $\epsilon_{t}:=(\epsilon_{1t},\ldots,\epsilon_{Nt})^{\top}$is allowed to be strong (e.g., with a factor structure) along the cross-sectional dimension and it has to be weak along the time dimension, as we only exploit the sample size in the time series dimension for the estimation of heterogeneous threshold effects. Specifically, we assume that $\epsilon_{t}$ follows an MA($\infty$) process as follows:
@@
Remark 1.
(Dependence between $X_{jt}$ and $U_{jt}$.) Note that Equation (3.1) is a fairly general setup, where almost all kinds of dependence structures between $X_{jt}$ and $U_{jt}$ can be fulfilled for the following simple reasoning. To see this, take $U_{jt}$ as a scalar random variable as an example. For any two continuous random variables $X$ and $U$ with the conditional cumulative distribution function of $U$ given $X$ as $F_{U|X}(\cdot|X)$. Let $\tilde{e}=F_{U|X}(U|X)$. Then by Lemma F.1 in Appendix F in Rio, (2017), $\tilde{e}$ is a uniformly distributed random variable on $[0,1]$ and is independent of $X$. Moreover, $U=F_{U|X}^{-1}(\tilde{e}|X).$            $\Box$
Next, we impose some conditions on the elements of $\eta_{t}.$
@@
The following assumption imposes some conditions on the temporal dependence for the processes $(X_{jt})_{t\in\mathbb{Z}}$ and $(\epsilon_{jt})_{t\in\mathbb{Z}}.$ Denote $\mathcal{C}_{j}$ as a set of search values for the time series $j$ containing at least the threshold values $c_{0j}$. If the true threshold is given, then $\mathcal{C}_{j}=\{c_{0j}\}$ is a single point in subsections 3.2 and A.3. We will specify $\mathcal{C}_{j}$ later in the theorems for the unknown case. Let $g_{j}(\cdot)$ be the density function of $X_{jt}.$ Recall $\mathcal{F}_{s}$ in (3.2), for $t\geq s+1,$ let $g_{j,t}(x|\mathcal{F}_{s}):=d\mathbb{P}(X_{jt}\leq x|\mathcal{F}_{s})/dx.$
@@
Assumption 3.2 $(i)$-$(ii)$ essentially assumes weak temporal dependence for the processes $(X_{jt})_{t\in\mathbb{Z}}$ and $(\epsilon_{jt})_{t\in\mathbb{Z}}$ for any $j\in\left[N\right]$. We allow strong cross-sectional dependence over the $j$ dimension. We shall also note that $\beta$ indicates the strength of the dependency structure of the error processes.
Furthermore, let $g_{j_{1},j_{2}}(\cdot,\cdot|\mathcal{F}_{t-1})$ be the conditional joint density of $X_{j_{1}t}$ and $X_{j_{2}t}$ given $\mathcal{F}_{t-1}.$ Finally, the following assumption imposes a condition on $g_{j_{1},j_{2}}(\cdot,\cdot|\mathcal{F}_{t-1}).$